Factor 55^4 into the product of prime factors by first factoring 55 and then raising each of its factors to the fourth power.
^-means the power to the number

Answer by ankor@dixie-net.com(22740)

Factor 55^4 into the product of prime factors by first factoring 55 and then raising each of its factors to the fourth power.
:
(5*11)^4 = 5^4 * 11^4 = 625 * 14641 = 9,150,625